OpenStudy (unklerhaukus):

\[y^{\prime\prime}+4y=\cos2x\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}+4y_c=0\]\[m^2+4m=0\]\[m(m+4)=0\]\[m=0,-4\]\[y_c=A+Be^{-4}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p=\cos2x\]\[y_p=C\cos2x+D\sin2x\]\[y_p^\prime=-2C\sin2x+2D\cos2x\]\[y_p^{\prime\prime}=-4C\cos2x-4D\sin2x\]

OpenStudy (unklerhaukus):

\[(-4C\cos2x-4D\sin2x)+4(C\cos2x+D\sin2x)=\cos2x\]\[2C\cos2x+2D\sin2x=\cos2x\]\[2D\sin2x=0;\qquad 2C\cos2x=\cos2x\]\[D=0;\qquad C=\frac{1}{2}\]\[y_p=\frac 12\cos2x\] \[y=y_c+y_p\] \[y(x)=A+Be^{-4}+\frac 12\cos2x\]

OpenStudy (unklerhaukus):

however the answer in the back of the book was \[ y (x)= \frac{1}4x\sin2x+A\cos2x+B\sin2x \] where did i go wrong/

OpenStudy (blockcolder):

You plugged in the wrong thing. You put y instead of y' two posts ago.

OpenStudy (unklerhaukus):

ah yes i got the substitution wrong

OpenStudy (unklerhaukus):

wait y' ?

OpenStudy (anonymous):

Well, I couldn't find the mistake yet..

OpenStudy (unklerhaukus):

was i ment to multiply the particular solution by x for some reason?

OpenStudy (anonymous):

Wait, I think you will get complex roots... \[y_c'' + 4y_c = 0\] \[m^2 + 4 = 0\] Isn't it??? \[m = \pm 2i\]

OpenStudy (unklerhaukus):

yeah i did make a mistak my my second lines of working, \[\cancel{m^2+4m=0}\] thanks for spotting that @waterineyes \[m^2+4=0\]

OpenStudy (anonymous):

Happy To Help... Welcome dear..

OpenStudy (unklerhaukus):

\[y_c=A\left(E\cos 2x+F\sin 2x\right)\]

OpenStudy (anonymous):

I think you can easily take as: \[y_c = Acos2x + B \sin2x\] Or not??

OpenStudy (unklerhaukus):

ah yes good idea

OpenStudy (unklerhaukus):

\[y^{\prime\prime}+4y=\cos2x\]\[y_c^{\prime\prime}+4y_c=0\]\[m^2+4=0\]\[m=\pm2i\]\[y_c=A\cos 2x+B\sin 2x\]

OpenStudy (anonymous):

Yes you are going right... Now solve for particular solution..

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p=\cos2x\]\[y_p=C\cos2x+D\sin2x\]\[y_p^\prime=-2C\sin2x+2D\cos2x\]\[y_p^{\prime\prime}=-4C\cos2x-4D\sin2x\]\[(-4C\cos2x-4D\sin2x)+4(C\cos2x+D\sin2x)=\cos2x\]\[0=\cos2x\] ??there must be a mistake

OpenStudy (anonymous):

How can this be possible?? I am in trouble too... See your answers first term is \(\frac{1}{4}xsin2x\) we have to get this term but how..

OpenStudy (anonymous):

Are you sure your question contains \(cos2x\) on Right Hand Side..??

OpenStudy (unklerhaukus):

yes the question in the book is \[y^{\prime\prime}-4y=\cos2x\]

OpenStudy (unklerhaukus):

oh no i see +≠-, what a mistake to make sorry everyone

OpenStudy (unklerhaukus):

hmm actually i have two copies of the book and they are different

OpenStudy (unklerhaukus):

the latest version the one i was working with , has made a printing error, the question should have been \[y^{\prime\prime}-4y=\cos2x\]

OpenStudy (unklerhaukus):

\[\text{---------}\]\[y^{\prime\prime}-4y=\cos2x\] \[y_c^{\prime\prime}-4y_c=0\] \[m^2-4=0\]\[m=\pm2\] \[y_c=Ae^{2x}+Be^{-2x}\] \[y_p^{\prime\prime}-4y_p=\cos2x\] \[y_p=C\cos2x+D\sin2x\]\[y_p^\prime=-2C\sin2x+2D\cos2x\]\[y_p^{\prime\prime}=-4C\cos2x-4D\sin2x\] \[(-4C\cos2x-4D\sin2x)+4(C\cos2x+D\sin2x)=\cos2x\]\[-8C\cos2x-8D\sin2x=\cos2x\]\[-8D\sin2x=0;\qquad -8C\cos2x=\cos2x\]\[D=0;\qquad C=-\frac{1}{8}\] \[y_p=-\frac 18\cos2x\] \[y=y_c+y_p\] \[y(x)=Ae^{2x}+Be^{-2x}-\frac 18\cos2x\]\[\text{---------}\]